OpenStudy (anonymous):

How many cubic yards of concrete are needed to build a concrete pipe 9 feet long and 9 inches thick if the interior diameter must be 54 inches? (round tot eh nearest hundredth)

OpenStudy (campbell_st):

if the interior diameter is 54 inches then the interior radius is 27 inches. the exterior radius will be 27 + 9 = 36 inches you need to find the area between the 2 circles...  but before calculating convert all your measurements to yards... since you need to find the volume in cubic yards. 27 inches = 0.75 yards 36 inches = 1 yard 9 ft = 3 yards find the area between the circles Area = Big Circle - Little Circle \[A = \pi \times 1^2 - \pi \times 0.75^2 \] when you know the area, to find the volume multiply by the length of the pipe \[V = A \times 3\] You will need to round your answer to 2 decimal places... hope this helps.

OpenStudy (anonymous):

THANKS!